# Elastic, Quasielastic, and Superelastic Electron Scattering from Thermal Lattice Distortions in Perfect Crystals

### v2

@@ V Elastic internal scattering and the Debye-Waller fraction
$\hat{\mathbf{u}}_{j}=\frac{1}{\sqrt{N}}\sum_{\mathbf{k},\lambda}\sqrt{\frac{\hbar}{2M\omega_{\mathbf{k}\lambda}}}\,\mathbf{e}_{\mathbf{k}\lambda}\left(a_{\mathbf{k}\lambda}e^{i\mathbf{k}\cdot\bm{\xi}^{(0)}_{j}}+a^{\dagger}_{\mathbf{k}\lambda}e^{-i\mathbf{k}\cdot\bm{\xi}^{(0)}_{j}}\right)$. (22)
The operator $e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}$ therefore contains all orders in the phonon operators and automatically generates elastic, quasielastic, and inelastic processes.
In particular, in a phonon number state basis the diagonal matrix elements are generically nonzero,
$\left\langle\{n\}\left|e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\right|\{n\}\right\rangle\neq 0,$ (23)
so that a scattering channel strictly elastic with respect to internal lattice excitations is present already at first order in $H_{\rm int}$. Upon thermal averaging (or in the ground state), this diagonal factor reduces to the Debye-Waller form
$\left\langle e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\right\rangle_{T}=\exp[-W(\mathbf{q},T)],$ (24)
which is the direct analog of the recoil-free (zero-phonon-line) fraction in Mössbauer spectroscopy and in neutron or x-ray scattering.
While much stronger coupling regimes can be treated nonperturbatively—for example within wave-on-wave (WoW) simulations [4] the present analysis shows that even within standard perturbation theory there exist elastic scattering processes that are absent when the center of mass is hidden or eliminated.
Relation to the Fröhlich truncation
In the original Fröhlich formulation, the exact factorization Eq. (11) is not employed. The lattice center-of-mass coordinate $\hat{\mathbf{R}}$ is not retained explicitly, which is equivalent to setting the recoil factor $e^{-i\mathbf{q}\cdot\hat{\mathbf{R}}}$ to unity and treating the lattice as a clamped reference frame. At the same time, the internal displacement exponential is truncated to linear order,
$e^{-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}}\rightarrow 1-i\mathbf{q}\cdot\hat{\mathbf{u}}_{j}.$ (25)
The zeroth-order term reproduces the static periodic potential responsible for Bloch band formation, while the linear term generates single-phonon processes.
Although this procedure is internally consistent and highly successful in the weak-coupling regime, it removes the background recoil zero mode by construction. As a result, even if the full exponential in $\hat{\mathbf{u}}_{j}$ were retained, the truncated theory would still lack an explicit degree of freedom capable of absorbing the total recoil momentum of the lattice. The standard Fröhlich Hamiltonian therefore describes momentum exchange with internal lattice modes, but the 0 mode, the lattice recoil as a whole, is absent.
Thermal averaging and the Debye-Waller fraction
@@
the familiar Debye–Waller result. Even at finite temperature a substantial fraction of momentum-transfer events leave the lattice in the same internal state.
Estimate of the elastic internal fraction at representative momentum transfers
@@
VI Experiments suggesting large elastic contributions
Experimental evidence for predominantly elastic momentum relaxation in metals at low temperature has accumulated for decades, particularly in the mesoscopic-transport and quantum-interference literature. In this section we show that the framework developed here resolves several longstanding disparities between theory and experiment regarding electronic transport. We review a range of experiments that demonstrate a clear separation between momentum relaxation and energy relaxation, consistent with elastic or quasi-elastic scattering dominating momentum randomization.
Hierarchy of momentum and energy relaxation times
@@ VI Experiments suggesting large elastic contributions
The consistent appearance of this hierarchy demonstrates that momentum relaxation and energy relaxation are distinct physical processes. While conventional language often associates momentum randomization with inelastic scattering, these experiments show that electronic momentum can be efficiently scrambled by a time-dependent environment without significant energy exchange. The lattice need not act as an energy sink in order to act as a momentum sink. This separation is precisely what is expected when electrons scatter elastically from a dynamically fluctuating lattice background.
Quantum interference and dephasing
@@
VII Discussion
The purpose of this section is twofold. First, we examine the consistency of the present framework with well-established experimental results, including the Wiedemann–Franz law, weak localization, and quantum oscillations. These considerations address potential concerns regarding the role of predominantly elastic momentum-relaxing channels. Second, we outline broader interpretive implications for Planckian transport, linear-in-$T$ resistivity, and the Mott–Ioffe–Regel crossover. The latter discussion is exploratory and is presented as a coherent physical interpretation rather than as a definitive claim.
VII.1 Implications for the Lorenz ratio
@@
VII.2 Weak localization
Weak localization (WL) provides a clear illustration of the central role of elastic backscattering in low-temperature metallic transport. The phenomenon arises from constructive interference between pairs of time-reversed electronic trajectories that revisit the same spatial region after multiple scattering events [6, 20]. The WL correction therefore requires that a substantial fraction of backscattering processes preserve phase coherence over the relevant timescale.
@@ VII.2 Weak localization
VII.3 Shubnikov–de Haas oscillations
Quantum oscillation phenomena such as the Shubnikov–de Haas and de Haas–van Alphen effects provide an independent probe of electronic scattering in metals at low temperature [29, 2]. These oscillations arise from the quantization of electronic orbits in a magnetic field and reflect coherent motion around the Fermi surface over many cyclotron periods.
The oscillation amplitude is controlled by the Dingle factor [8],
$\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}}\right),$ (37)
where $\tau_{q}$ is the quantum lifetime associated with Landau-level broadening. This lifetime is sensitive to momentum-randomizing processes that broaden Landau levels, while sufficiently strong inelastic energy exchange suppresses the coherence required to complete cyclotron motion. Scattering events that transfer energies comparable to $\hbar\omega_{c}$ disrupt the phase coherence of the orbit.
The observation of well-defined quantum oscillations over broad temperature ranges and down to relatively low magnetic fields therefore indicates that electrons can undergo many momentum-deflecting collisions while maintaining phase coherence over multiple cyclotron periods. As in weak localization, the temperature dependence of the oscillation amplitude reflects phase smearing due to temporal fluctuations of the scattering environment. These considerations are consistent with a regime in which momentum randomization can occur without rapid or large energy exchange on electronic scales.
VII.4 Planckian diffusion and linear-in-$T$ resistivity
@@
up to factors of order unity. From this viewpoint, the Planckian timescale may be understood not as a fundamental bound on inelastic scattering, but as an emergent parametrization of quantum-limited diffusion combined with equilibrium thermodynamics. It characterizes the rate at which charge spreads spatially rather than the rate at which energy must be irreversibly transferred to internal degrees of freedom.
VIII Summary
In the conventional clamped-lattice formulation of electron–phonon scattering, the lattice is treated as a fixed reference frame and its translational zero mode does not appear explicitly among the dynamical variables. Within that representation, momentum transfer is naturally described in terms of internal phonon excitations. When the lattice center-of-mass degree of freedom is retained explicitly, an additional elastic momentum-transfer channel becomes manifest. In a finite crystal, pseudomomentum can be exchanged between an electron and the lattice background while total mechanical momentum is conserved through recoil of the lattice zero mode. Such processes do not require phonon creation and may contribute significantly to momentum relaxation at low and intermediate temperatures.
Making this channel explicit requires adopting a formulation in which (i) the lattice is treated as a finite, mechanically isolated object so that its center-of-mass motion remains dynamical, (ii) discrete translational symmetry is implemented symmetrically, with foreground and background pseudomomenta treated on equal footing, and (iii) the electron–lattice interaction is retained in its full density–density form. These steps do not alter the microscopic Hamiltonian; rather, they retain degrees of freedom that are commonly suppressed by choice of reference frame. When treated in this manner, elastic momentum exchange between electron and lattice background appears naturally within the same theoretical structure.
In this framework, pseudomomentum is redistributed between the electron and the lattice background, while total mechanical momentum is conserved through recoil of the translational zero mode. Momentum relaxation need not coincide with phonon creation, and energy dissipation need not set the transport rate. This separation of roles clarifies how efficient momentum randomization can coexist with comparatively slow energy relaxation.
Restoring explicit center-of-mass dynamics provides a more complete description of momentum bookkeeping in finite crystals. It offers a unified interpretation of experimental observations that indicate robust elastic scattering in clean materials, including weak localization, quantum oscillations, and ultrasonic attenuation. More broadly, the results suggest that in certain regimes, including some materials exhibiting linear-in-$T$ resistivity, momentum relaxation may be governed primarily by elastic scattering in a time-dependent lattice background [5], while inelastic processes play a secondary role in establishing global thermal equilibrium.
In a perfectly clean, isolated, translation-invariant crystal, total momentum is conserved exactly. Such a system may exhibit internal equilibration and decay of electronic current, but it cannot exhibit finite DC resistivity in the strict steady-state sense. True resistivity requires a mechanism for removing total momentum from the system.
IX Acknowledgments

### v3

@@ VII.3 Recoil-dominated channel
$\Delta E_{\rm recoil}=\frac{\hbar^{2}}{2M_{\rm tot}}\left[(\mathbf{K}-\mathbf{q})^{2}-\mathbf{K}^{2}\right]=\frac{(\hbar q)^{2}}{2M_{\rm tot}}-\frac{\hbar^{2}}{M_{\rm tot}}\mathbf{K}\!\cdot\!\mathbf{q},$ (77)
which scales as $1/M_{\rm tot}$ and therefore vanishes in the thermodynamic limit. To leading order in $1/N$, the scattering is thus elastic with respect to internal lattice excitations.
The internal factor
$\langle\Psi_{f}^{\rm bg}|\hat{\rho}^{\rm int}_{-\mathbf{q}}|\Psi_{i}^{\rm bg}\rangle$ (78)
remains finite even when no phonon quanta are created or annihilated. For a thermally fluctuating background its diagonal contribution reduces, upon averaging, to the Debye-Waller factor $e^{-W(\mathbf{q},T)}.$
VII.4 Quasielastic and superelastic processes
More generally, $\hat{\rho}^{\rm int}_{-\mathbf{q}}$ contains all orders in internal lattice operators. Transitions therefore exist in which the center-of-mass absorbs the entire pseudomomentum transfer, while the internal modes undergo small changes that adjust the energy balance. Such processes are quasielastic:
$\epsilon_{n^{\prime}\mathbf{k}^{\prime}}\approx\epsilon_{n\mathbf{k}},\qquad|\Delta E_{\rm int}|\ll\epsilon_{n\mathbf{k}},$ (79)
with the exact kinematic constraint always enforced by center-of-mass recoil.
At finite temperature, superelastic channels occur when pre-existing internal excitations are partially annihilated, transferring energy to the electron while the net pseudomomentum remains balanced by the center-of-mass degree of freedom.
VII.5 Elastic and quasi-elastic processes beyond the Fröhlich term
Consider the electron-lattice interaction in its full density-density form
$H_{\rm int}=\sum_{q}V(q)\,\rho_{e}(-q)\,\rho_{\rm bg}(q),\qquad\rho_{e}(-q)=\sum_{k}c^{\dagger}_{k+q}c_{k}.$ (80)
The background density may be written as
$\rho_{\rm bg}(q)$ $=\sum_{j}e^{-iq(R_{j}+u_{j})}$ (81)
$=\sum_{j}e^{-iqR_{j}}\left(1-iq\,u_{j}-\frac{q^{2}}{2}u_{j}^{2}+\cdots\right).$
where $R_{j}$ are equilibrium positions and $u_{j}$ displacements.
The zeroth-order term gives
$H^{(0)}=\sum_{k,q}V(q)\,c^{\dagger}_{k+q}c_{k}\sum_{j}e^{-iqR_{j}},$ (82)
which contains no phonon operators and is therefore purely elastic. In a perfect periodic crystal this reduces to reciprocal lattice scattering ($q=G$), i.e. Bragg processes. It is responsible for the static band structure. As Kohn noted[20], there is plenty of elastic deflection going into the construction of bands, which gets masked in making Bloch waves. Figure 5 makes this clear.
Figure 5: Interrupted Bloch wave Both panels feature a circular absorbing zone seen clearly, as the real part of a plain wave (left) and of a Bloch wave (right) enter from above. A shadow persists and slowly fills by diffraction on the left, plane wave case, but strong scattering by the triangular lattice present on the right quickly fills in the void. This scattering, which is obscured by periodic coherence in the unobstructed Bloch waves, is plainly revealed as the void fills. The point is that the strong unprotected scattering in the shadow region cannot be strongly inelastic, or else passive barriers in periodic lattices would cause heating downstream.
The linear term
$H^{(1)}=-i\sum_{k,q}q\,V(q)\,c^{\dagger}_{k+q}c_{k}\sum_{j}e^{-iqR_{j}}u_{j},$ (83)
yields, upon expanding $u_{j}$ in phonon operators, the familiar Fröhlich interaction in which each term carries a definite phonon pseudomomentum.
At quadratic order one finds
$H^{(2)}=-\frac{1}{2}\sum_{k,q}q^{2}V(q)\,c^{\dagger}_{k+q}c_{k}\sum_{j}e^{-iqR_{j}}u_{j}^{2},$ (84)
and writing
$u_{j}=\frac{1}{\sqrt{N}}\sum_{Q}U_{Q}\,e^{iQR_{j}}(a_{Q}+a^{\dagger}_{-Q}),$ (85)
one obtains terms of the form
$\begin{split}H^{(2)}\sim\sum_{k,q}\sum_{Q,Q^{\prime}}c^{\dagger}_{k+q}c_{k}\,(a_{Q}+a^{\dagger}_{-Q})(a_{Q^{\prime}}+a^{\dagger}_{-Q^{\prime}})\\ \times\sum_{j}e^{-i(q-Q-Q^{\prime})R_{j}}.\end{split}$ (86)
Among these are phonon-diagonal contributions with $Q^{\prime}=-Q$,
$H^{(2)}_{\rm el}\sim\sum_{k,q,Q}c^{\dagger}_{k+q}c_{k}\,\bigl(2a_{Q}^{\dagger}a_{Q}+1\bigr)\sum_{j}e^{-iqR_{j}},$ (87)
which have nonzero matrix elements
$\langle\nu|H^{(2)}_{\rm el}|\nu\rangle\neq 0.$ (88)
These describe elastic electron scattering with $\nu\to\nu$, and more generally quasi-elastic processes when $\nu\to\nu^{\prime}$ with small energy change.
Thus, once the interaction is treated beyond the linear Fröhlich approximation, elastic and quasi-elastic channels appear already at first order in the density-density interaction. In such processes $\Delta K_{\rm ph}=0$, while in general $\Delta K_{e}\neq Gm$, so the standard selection rule
$\Delta K_{e}+\Delta K_{\rm ph}=Gm$ (89)
is not satisfied within the zero-mode-removed theory.
This is consistent with the observation, emphasized by Kohn, that the electronic band structure itself arises from coherent elastic scattering from the periodic lattice potential. Elastic scattering is therefore not an exceptional process but the dominant one that defines the electronic states. The restriction to inelastic, one-phonon processes is a feature of the linearized phonon representation, not of the underlying dynamics.
The appearance of elastic and quasi-elastic channels thus exposes the limitations of the reduced description. Once the lattice center-of-mass degree of freedom is omitted, neither mechanical momentum nor pseudomomentum is properly accounted for within the phonon subspace alone. The full problem requires a dynamical background capable of carrying the compensating momentum.
VII.6 Thermal averaging and the Debye-Waller fraction
@@
the familiar Debye-Waller result. Even at finite temperature a substantial fraction of momentum-transfer events leave the lattice in the same internal state.
VII.7 Estimate of the elastic internal fraction at representative momentum transfers
@@
We can motivate this result as follows: For a Cu atom modeled as an Einstein oscillator, a sudden momentum transfer $\Delta p=\hbar q$ leaves the oscillator in its ground state with probability
$P_{0\to 0}=\left|\langle 0|e^{iqx}|0\rangle\right|^{2}=\exp\!\left(-\frac{E_{R}}{\hbar\omega_{E}}\right),\qquad E_{R}=\frac{(\hbar q)^{2}}{2M}.$ (98)
Taking $q=2k_{F}$ for copper, with $k_{F}\approx 1.36\,\mathrm{\AA }^{-1}$ and $M\approx 63.5\,u$, gives $E_{R}\approx 0.24\,\mathrm{meV}$. For a typical local vibrational energy $\hbar\omega_{E}\sim 20$-$30\,\mathrm{meV}$, one finds $P_{0\to 0}\approx 0.99$.
VIII Experiments suggesting large elastic contributions
We have already touched on pure wire resistivity and persistent currents above, as allowing a more satisfactory explanation of the experiments through inclusion of the zero modes and the expansion to elastic and quasielastic channels that they enable.
Other experimental evidence suggesting predominantly elastic momentum relaxation in metals at low temperature has accumulated for decades, particularly in the mesoscopic-transport and quantum-interference literature. In this section we show that the framework developed here can resolve several longstanding disparities between theory and experiment regarding electronic transport. We review a range of experiments that demonstrate a clear separation between momentum relaxation and energy relaxation, consistent with elastic or quasi-elastic scattering dominating momentum randomization.
VIII.1 Hierarchy of momentum and energy relaxation times
@@
The consistent appearance of this hierarchy demonstrates that momentum relaxation and energy relaxation are distinct physical processes. These experiments show that electronic momentum can be efficiently scrambled by a time-dependent environment without significant energy exchange. In other words, the lattice need not act as an energy sink in order to act as a momentum sink. This separation is precisely what is expected when electrons scatter quasi-elastically from a dynamically fluctuating lattice background.
VIII.2 Quantum interference and dephasing
@@ VIII.3 Hot-electron relaxation
Johnson-noise thermometry provides a direct probe of electron–lattice energy exchange. Experiments consistently show that the electron temperature relaxes on timescales much longer than those associated with momentum relaxation inferred from transport[11]. This separation, $\tau_{p}\ll\tau_{E}$, is difficult to reconcile with a picture in which momentum relaxation is intrinsically inelastic. It is, however, naturally explained if momentum is transferred predominantly through elastic or quasielastic processes, while energy relaxation proceeds through slower coupling to internal lattice degrees of freedom.
VIII.4 Shubnikov-de Haas oscillations
Quantum oscillation phenomena such as the Shubnikov-de Haas and de Haas-van Alphen effects provide a precise probe of electronic coherence and scattering in metals at low temperature [33, 3]. These oscillations arise from Landau quantization and require that electrons execute many cyclotron orbits without loss of phase coherence.
Within the conventional Lifshitz-Kosevich (LK) framework, the oscillation amplitude may be written as
$A(T,B)\propto\frac{X}{\sinh X}\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}}\right),\qquad X=\frac{2\pi^{2}k_{B}T}{\hbar\omega_{c}},$ (100)
where the factor $X/\sinh X$ describes thermal smearing of the Fermi surface, while the Dingle factor
$\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}}\right)$ (101)
encodes Landau-level broadening through the quantum lifetime $\tau_{q}$. In standard treatments, $\tau_{q}$ is associated with inelastic or quasi-inelastic processes, and is often assumed to become weakly temperature dependent at low $T$.
From the present viewpoint, this interpretation is unnecessarily restrictive. When the lattice center-of-mass degree of freedom is retained, momentum relaxation need not proceed through phonon creation or annihilation. Instead, the electron can exchange momentum elastically with the lattice background via recoil of the zero mode, while total mechanical momentum is conserved. These processes are elastic in energy, but nonetheless randomize the electronic phase.
As a result, the Dingle factor should be understood more generally as encoding phase randomization rather than strictly inelastic scattering. Elastic recoil processes provide a continuous channel for Landau-level broadening, even in the absence of phonon excitation. The temperature dependence then arises not from phonon occupation factors, but from the growth of available recoil phase space in the thermally fluctuating background.
This leads naturally to a temperature-dependent quantum lifetime of the form
$\frac{1}{\tau_{q}(T)}\sim T^{\alpha},$ (102)
with $\alpha\sim 1$ in a Planckian regime, implying
$\exp\!\left(-\frac{\pi}{\omega_{c}\tau_{q}(T)}\right)\;\sim\;\exp\!\left(-\mathrm{const}\,\frac{T}{B}\right).$ (103)
Thus, the damping of quantum oscillations can remain strongly temperature dependent down to low $T$ without invoking inelastic scattering.
The persistence of Shubnikov-de Haas oscillations over wide temperature ranges therefore indicates that substantial momentum randomization can occur through predominantly elastic processes, while phase coherence is degraded only gradually. This resolves the apparent tension between strong scattering and long-lived cyclotron coherence, and places quantum oscillation damping on the same footing as weak localization: both probe phase randomization in a dynamically fluctuating, but not necessarily energy-relaxing, environment.
This interpretation is consistent with experimental observations of a marked temperature dependence of the Dingle factor, often extending to low $T$ with approximately power-law or linear-in-$T$ behavior, rather than saturation. In contrast to conventional theory, which attributes Landau-level broadening primarily to inelastic processes and therefore expects weak temperature dependence at low $T$, the present framework predicts a continued $T$-dependent damping arising from elastic recoil-driven phase randomization.
IX Discussion
Here we examine the consistency of the present framework with well-established experimental results, including the Wiedemann–Franz law, weak localization, and quantum oscillations.
Second, we outline broader interpretive implications for Planckian transport, linear-in-$T$ resistivity, and the Mott–Ioffe–Regel crossover. The latter discussion is exploratory and is presented as a coherent physical interpretation rather than as a definitive claim.
IX.1 Implications for the Lorenz ratio
@@
IX.2 Weak localization
Weak localization (WL) provides a clear illustration of the central role of elastic backscattering in low-temperature metallic transport. The phenomenon arises from constructive interference between pairs of time-reversed electronic trajectories that revisit the same spatial region after multiple scattering events [7, 23]. It is often the case that the WL correction is important only with defects causing elastic backscattering. However the electrons still have to arrive back intact, so to speak (i.e. in phase), so other inelastic events cannot have decohered the electron on the way out and back. This is a tall order if any “natural” deflections are inelastic. elcrtherefore requires that a substantial fraction of backscattering processes preserve phase coherence over the relevant timescale.
@@ IX.2 Weak localization
IX.3 Planckian diffusion and linear-in-$T$ resistivity
@@
up to factors of order unity. From this viewpoint, the Planckian timescale may be understood not as a fundamental bound on inelastic scattering, but as an emergent parametrization of quantum-limited diffusion combined with equilibrium thermodynamics. It characterizes the rate at which charge spreads spatially rather than the rate at which energy must be irreversibly transferred to internal degrees of freedom, and thus fits nicely into the main elastic and quasielastic theme of this paper.
IX.4 Coulomb drag
We began discussing Coulomb drag in section II. It is an ideal case in which the zero recoil mode, missing in the literature on coulomb drag, is likely the agent of the power law seen in experiments. Theoretical analyses of one-dimensional transport consistently emphasize the strong kinematic constraints imposed by simultaneous energy and momentum conservation, leading to the conclusion that relaxation processes should be strongly suppressed at low temperature [26, 18] . Experimentally, however, relaxation and drag in quantum wires are often observed to exhibit power-law temperature dependence rather than strong suppression. This discrepancy is typically attributed to additional mechanisms such as disorder, band curvature, or coupling to external degrees of freedom that relax the kinematic constraints. This will be the topic of a future publication.
IX.5 Phonon drag
In conventional transport theory, the phonon-drag contribution to the thermopower is attributed to momentum transfer from a nonequilibrium phonon population to the electrons, and is therefore tied to phonon creation and annihilation. This reflects a formulation in which the lattice center-of-mass (zero mode) is removed and momentum exchange is forced into the phonon sector.
When the zero mode is retained, momentum transfer is no longer tied to phonon number. The electron couples to the lattice deformation field, and momentum can be exchanged with the lattice as a whole through recoil, without requiring phonon excitation. The conventional phonon-drag contribution therefore represents only a subset of a more general lattice-drag response.
Accordingly, the thermopower should be written as
$S=S_{\rm diff}+S_{\rm lat},\qquad S_{\rm lat}=S_{\rm el}+S_{\rm inel},$ (109)
where $S_{\rm inel}$ is the usual phonon-drag term and $S_{\rm el}$ arises from elastic or quasielastic momentum transfer via lattice recoil. In regimes where scattering is predominantly elastic, $S_{\rm el}$ need not be small and may dominate the drag response.
X Summary
We have revisited electron-lattice scattering starting from the standard microscopic Hamiltonian, but retaining a degree of freedom that is nearly always removed: the translational center-of-mass (zero mode) of the lattice. In conventional formulations, this mode is eliminated (e.g., through Born-von Kármán boundary conditions), leaving momentum transfer to be carried entirely by internal phonon excitations. Together with the form of the first order Fröhlich expansion, changes in electronic momentum are tied to phonon creation or annihilation, and scattering is treated as intrinsically inelastic.
When the lattice center of mass is retained explicitly, and the density-density form of the interaction is retained and not expanded, this picture changes qualitatively. Total mechanical momentum is conserved by recoil of the lattice as a whole, while pseudomomentum is redistributed between electron and background. A broad class of elastic and quasielastic momentum-transfer processes then becomes available, in which electronic momentum changes are balanced by the zero mode without requiring phonon excitation.
This enlargement of the allowed phase space follows directly from the exact symmetries of the Hamiltonian. No new interaction has been introduced. Retaining the full density-density interaction reveals that elastic matrix elements arise already at the level of the state-to-state transition amplitude $\langle f|H_{\rm int}|i\rangle$, without invoking higher-order or multiphonon processes. The familiar phonon-mediated processes remain, but represent a restricted subset of a much larger set of allowed transitions.
The resulting kinematics has broad implications. In one-dimensional systems, the need for thermally activated three-particle processes is removed, and power-law relaxation follows naturally. More generally, momentum relaxation need not be tied to energy dissipation: momentum can randomize rapidly while energy relaxes on much longer timescales. This resolves a long-standing tension between theory and experiment in clean metals, including observations from weak localization, quantum oscillations, hot-electron transport, and Coulomb drag.
We emphasize that these conclusions are not model-dependent. Once the full momentum conservation law of the isolated electron-lattice system is enforced, the existence of elastic and quasielastic scattering channels is unavoidable. The restriction to phonon-mediated inelastic scattering is therefore not a fundamental feature of the physics, but a consequence of having discarded the lattice center-of-mass degree of freedom.
XI Acknowledgments